OpenStudy (tal_tal):

The following data show the weight, in ounces, of 10 different bags of candy: 10 3 7 3 4 21 6 10 1 2 4 After removing the outlier, what does the mean absolute deviation of this data set represent? On average, the weight of a bag of candy varies 3.2 ounces from the mean of 4 ounces. On average, the weight of a bag of candy varies 2.6 ounces from the mean of 5 ounces. On average, the weight of a bag of candy varies 3.2 ounces from the mean of 5 ounces. On average, the weight of a bag of candy varies 2.6 ounces from the mean of 4 ounces.

OpenStudy (anonymous):

@bibby oh here better get her kik too

OpenStudy (tal_tal):

@dako87

rebeccaxhawaii (rebeccaxhawaii):

this isnt facebook ^.^ @sheyennepowell_

OpenStudy (anonymous):

I'm saying that because the bibby guy collects kiks from girls on here @rebeccaxhawaii

OpenStudy (dako87):

arent yall here to help lol

OpenStudy (welshfella):

what value is the outlier? any idea?

OpenStudy (tal_tal):

2?

OpenStudy (tal_tal):

I'm bad at these XD

OpenStudy (anonymous):

@bibby better collect this ones kik too

OpenStudy (welshfella):

No and ou;ier is a value thats a large amount less or greater than the other values

OpenStudy (kropot72):

There are 11 values of weight listed, although the question states that there are 10 different bags. It seems that there is an error somewhere.

OpenStudy (tal_tal):

@sheyennepowell_ why? xD

OpenStudy (tal_tal):

Ummm 21?

OpenStudy (welshfella):

- the question says the outlier is first removed

OpenStudy (anonymous):

because hes a rapist

OpenStudy (tal_tal):

? um okay

OpenStudy (welshfella):

yea 21 is correct can you work out the mean of the other 10 values?

OpenStudy (tal_tal):

5?

OpenStudy (anonymous):

@rebeccaxhawaii

OpenStudy (anonymous):

thats @bibby

OpenStudy (welshfella):

right 50 / 10 = 5 so do you know how to calculate the MAD? Mean Absolute Deviation

OpenStudy (tal_tal):

umm not really can you explain it

OpenStudy (welshfella):

OK Subtract each of the values in the list from the mean 5. If the result is negative change it to positive ( that's the absolute value) so for 10 we have 5-10 = -5 = 5 then for 3 we have 5 - 3 = 2 and so on ok?

OpenStudy (tal_tal):

oooh okay

OpenStudy (welshfella):

so do that and get the total of all 10 results what do you get?

OpenStudy (tal_tal):

well I have a quick question when I subtract 5 - 3 it's 2 but when it's 5 - 7 would that be -2?

OpenStudy (welshfella):

yes but the -2 becomes 2

OpenStudy (welshfella):

all values must be positive

OpenStudy (tal_tal):

oh okay thank you I will write my results

OpenStudy (tal_tal):

5-7=2 5-3=2 5-4=1 5-6=1 5-10=5 5-1=4 5-2=3 5-4=1

OpenStudy (tal_tal):

and then the two you already did lol

OpenStudy (welshfella):

yes the total of all 10 is 26 Now to find the MAD you divide this by 10

OpenStudy (welshfella):

MAD = 2.6 so now you can see which is the correct option

OpenStudy (tal_tal):

D?

OpenStudy (tal_tal):

I mean B

OpenStudy (welshfella):

Yes B is correct The MAD is a measure of spread of the data around the mean.

OpenStudy (tal_tal):

Thank you!

OpenStudy (welshfella):

yw

OpenStudy (tal_tal):

can you help me with another one?

OpenStudy (welshfella):

the formula for MAD is 

OpenStudy (welshfella):

where x with the bar above it is the mean and n is the number of values. The Big E means 'the sum of'.

OpenStudy (welshfella):

Sorry gtg right now

OpenStudy (tal_tal):

Okay! bye!